Phosphorous pentachloride is produced by the reaction between phosphine and chlorine gas. What mass of chlorine gas would be required to produced 0.015 kg of phosphorous pentachloride
1
Expert's answer
2019-04-01T07:47:46-0400

PH3+4 Cl2 = PCl5 + 3 HCl

 Mass of 1 mol of PCl5 = 31+5*35.5 = 208,5 g

0,015 kg /0.2085 kg/mol = 0,071942 mol

Mass of clorine = 4 *( 35,5*2) *0,071942 mol = 20.43 g